a student is asked to prepare 250cm3 of 0.1mol/dm3 sodium carbonate.he weighed 2.65g of sodium carbonate decahydrate (Na2CO3 10H2O) instead of anhydrous sodium carbonate( Na2CO3).Calculate the concentration of the solution in mol/dm3.
1
Expert's answer
2019-09-12T06:00:22-0400

Cm=n/V

n= 2.65/286= 0.00926 mol

V= 250 sm3 =0.25 dm3

Cm= 0.00926 mol/0.25 dm3 = 0.0370629371 mol/dm3